OpenStudy (anonymous):

Find the derivative, y=(sqrt(x))* (x+6)

OpenStudy (anonymous):

you could write it as \[y=x^{\frac{3}{2}}+6x^{\frac{1}{2}}\] and then use the almighty power rule

OpenStudy (anonymous):

whats the almighty power rule? lol

OpenStudy (anonymous):

\[\frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

Im still a little confused lol

OpenStudy (dls):

I think the question is \[\large \frac{d}{dx} \sqrt{x} \times (x+6)\] @satellite73 Use almighty product rule.

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i mean why use the product rule when you can multiply out and use the power rule?

OpenStudy (shamil98):

sqrt x = x^1/2

OpenStudy (shamil98):

use the product rule .

OpenStudy (dls):

okay i see what u did there

OpenStudy (shamil98):

f(x) = x^1/2 f'(x) = x^-1/2 g(x) = x+6 g'(x) = 1 just input those values in the product rule. .-.

OpenStudy (shamil98):

i dont get it what did satellite do o.o

OpenStudy (dls):

\[\Huge \sqrt{x} \times (x+1)= x \sqrt x+\sqrt{x} = x^\frac{3}{2} + x^\frac{1}{2}\] -.-

OpenStudy (shamil98):

oh derp

OpenStudy (dls):

:P

OpenStudy (dls):

it is x+6 btw just took an example

OpenStudy (anonymous):

okay thanks!